# Channel-Imposed Fusion: A Simple yet Effective Method for Medical Time Series Classification

### v1

@@ Abstract
The automatic classification of medical time series signals, such as electroencephalogram (EEG) and electrocardiogram (ECG), plays a pivotal role in clinical decision support and early detection of diseases. Although Transformer based models have achieved notable performance by implicitly modeling temporal dependencies through self-attention mechanisms, their inherently complex architectures and opaque reasoning processes undermine their trustworthiness in high stakes clinical settings. In response to these limitations, this study shifts focus toward a modeling paradigm that emphasizes structural transparency, aligning more closely with the intrinsic characteristics of medical data. We propose a novel method, Channel Imposed Fusion (CIF), which enhances the signal-to-noise ratio through cross-channel information fusion, effectively reduces redundancy, and improves classification performance. Furthermore, we integrate CIF with the Temporal Convolutional Network (TCN), known for its structural simplicity and controllable receptive field, to construct an efficient and explicit classification framework. Experimental results on multiple publicly available EEG and ECG datasets demonstrate that the proposed method not only outperforms existing state-of-the-art (SOTA) approaches in terms of various classification metrics, but also significantly enhances the transparency of the classification process, offering a novel perspective for medical time series classification.
@@ 1 Introduction
Medical time series (MedTS) data, such as electroencephalogram (EEG) and electrocardiogram (ECG) signals, are widely used in clinical settings to monitor patient health and play a crucial role in diagnosing neurological and cardiovascular diseases Arif et al. (2024); Xiao et al. (2023); Zhu et al. (2025). Accurate classification of these signals enables early anomaly detection, personalized treatment, and optimized therapy planning, ultimately improving patient outcomes and healthcare efficiency Liu et al. (2024a); Tian et al. (2023). With advances in deep learning, CNN-based models like EEGNet Lawhern et al. (2018) can automatically extract informative features from raw signals, significantly improving classification performance.
In recent years, Transformer models Vaswani et al. (2017), powered by the self-attention mechanism Bahdanau et al. (2014), have achieved remarkable progress in time-series modeling, particularly in capturing long-range dependencies and global contextual information Liu et al. (2021); Zhou et al. (2021). By converting sequential data into high-dimensional token embeddings, Transformers are capable of modeling complex temporal dependencies implicitly. However, despite their strong performance across a range of time-series tasks, several limitations have emerged when applying Transformer-based architectures to medical time-series classification, particularly in the following aspects: (1) Conflict Between Implicit Temporal Modeling and Explicit Physiological Mechanisms. Conventional Transformer architectures employ self attention mechanisms to implicitly represent temporal dependencies . However, essential physiological dynamics in medical signals, including conduction delays between ECG leads and synchronization of EEG rhythms, require explicit structural representations. Although methods such as Medformer Wang et al. (2024) have adopted strategies involving multiple scales of embedding, their attention distributions still fail to directly correspond to the underlying biological processes. This leads to fundamental limitations in model interpretability. (2) Mismatch Between Medical Domain Knowledge and Architectures Designed for General Purposes. Mainstream time series modeling methods such as Autoformer Wu et al. (2021), Informer Zhou et al. (2021), and FEDformer Zhou et al. (2022) have achieved strong performance in general domains like weather forecasting and finance. However, their architectural designs fail to account for the unique structural characteristics of medical time series. Medical signals often exhibit high inter-channel redundancy, spatial dependencies among physiological leads, and distinct local morphological patterns. These important features are easily weakened or overlooked in the high-dimensional and implicit modeling processes of conventional Transformer architectures.
To address these limitations, we rethink the modeling paradigm from the perspective of data representation and break away from the conventional approach of implicitly discovering unknown temporal relationships. Unlike existing methods that rely on implicit learning, we advocate for explicitly modeling temporal relationships that have clear physical or physiological foundations. Based on this principle, we propose the Channel-Imposed Fusion (CIF) method, which encodes prior causal structures directly into the feature representation. Specifically, CIF constructs new features through a linear combination of the form , where and represent signals from two different channels, and and are predefined coefficients derived from domain knowledge. By embedding such structured prior knowledge into the model, CIF not only enhances the ability to capture critical pathological patterns but also significantly improves robustness and interpretability when dealing with noisy or irregular medical time-series data. This approach provides a more intuitive and effective framework for medical time-series analysis, contributing to the development of reliable clinical decision support systems.
Based on the design philosophy of the CIF method, we aim for the model to maintain high efficiency while ensuring good interpretability and simplifying the structure as much as possible. To this end, we propose a simple yet effective Hidden-layer Mixed Bidirectional Temporal Convolutional Network (HM-BiTCN) architecture. This architecture utilizes a hierarchical causal convolution module, which explicitly encodes the coupling relationships between channels and temporal dynamics, effectively avoiding the issues of poor interpretability and structural complexity caused by the implicit modeling of temporal dependencies in Transformer-based methods. Moreover, it is worth noting that the CIF method is not limited to the HM-BiTCN architecture; it has high portability and can be seamlessly integrated with existing Transformer architectures to enhance the model’s adaptability and interpretability in medical time-series data. Our contributions are as follows:
(1) Proposing the Channel-Imposed Fusion (CIF) method: We propose the CIF method to explicitly model the relationships between channel dimensions in medical time-series data, particularly for signals with clear physiological structures, such as EEG and ECG. (2) Designing the HM-BiTCN model based on CIF: We leverage the CIF method to construct the Hidden-layer Mixed Bidirectional Temporal Convolutional Network (HM-BiTCN), which explicitly models channel coupling and temporal dependencies. By adopting this TCN-based approach, HMBiTCN breaks the dominance of transformer models (SOTA) in MedTS classification and surpasses them across multiple datasets. (3) Providing methodological transferability: The CIF method is not only applicable to the HM-BiTCN model but can also be seamlessly integrated into existing Transformer architectures.
@@ 2 Related Work
Medical Time Series Classification. Medical time series analysis diverges fundamentally from general time series forecasting Wu et al. (2022); Lu et al. (2024) by prioritizing pathological signature decoding over temporal extrapolation, with modalities like EEG Tang et al. (2021); Yang et al. (2023); Qu et al. (2020), ECG Xiao et al. (2023); Wang et al. (2023); Kiyasseh et al. (2021), and EMG [Xiong et al. (2021); Dai et al. (2022)] encoding distinct clinical semantics. Early methods were dominated by compact CNNs such as EEGNet Lawhern et al. (2018), which employs depthwise separable convolutions to efficiently extract spatio–temporal features while providing preliminary interpretability via feature-map visualization. Subsequently, temporal convolutional networks (TCNs) Bai et al. (2018); Lin et al. (2019) leveraging dilated causal convolutions achieved parallelizable computation and extended receptive fields, surpassing LSTM-based approaches Zhou et al. (2016); Shen and Lee (2016); Hochreiter and Schmidhuber (1997) on multiple medical signal classification benchmarks. Hybrid architectures such as EEG-Conformer Song et al. (2022) combined convolutional front-ends with Transformer self-attention to capture both local and global dependencies and enabled attention-based interpretability. More recently, fine-grained Transformer models such as Medformer Wang et al. (2024) introduced cross-channel tokenization and dual-stage self-attention, setting new SOTA accuracy on several public datasets. The latest MedGNN Fan et al. (2025) further augments attention mechanisms with multi-resolution graph learning to jointly model spatial multi-scale channel dependencies and temporal dynamics. Although Transformer models have significantly improved classification accuracy, their complexity, especially due to deep architectures and multi-layer attention mechanisms, results in reduced interpretability. In clinical applications, where model transparency and reliability are essential, it is imperative to strike a balance between high accuracy and explainability.
@@
Implicit modeling strategies in Transformer-based temporal methods. In time series analysis, Transformer-based models implicitly learn complex dependencies through diversified architectural designs: the vanilla Transformer Vaswani et al. (2017) first introduced multi-head self-attention and sinusoidal positional encoding to model temporal correlations globally; Informer Zhou et al. (2021) employs ProbSparse attention to select key time steps and compress sequence length, thereby reducing the computational cost of long-range dependencies; Reformer Kitaev et al. (2019) incorporates Locality-Sensitive Hashing (LSH) to reduce attention complexity to , making it suitable for ultra-long sequences; Autoformer Wu et al. (2021) proposes an Auto-Correlation mechanism that aggregates periodic subsequences to enhance the implicit capture of cyclic patterns; FEDformer Zhou et al. (2022) performs seasonal–trend decomposition in the frequency domain and uses compressed Fourier coefficients to enable cross-frequency attention interactions; Crossformer Zhang and Yan (2022) designs a two-stage attention mechanism across time and feature dimensions to implicitly fuse multivariate spatiotemporal couplings; iTransformer Liu et al. (2024b) innovatively treats time steps as channel dimensions and applies standard attention to implicitly learn nonlinear inter-variable relationships; PatchTST Nie et al. (2023) segments continuous time steps into patch-based tokens and uses a combination of local and global attention to capture multi-scale temporal patterns; Medformer Wang et al. (2024) introduces multi-granularity patch embeddings and cross-channel attention for medical signals, implicitly modeling the heterogeneous couplings of physiological metrics; and MedGNN Fan et al. (2025) combines graph attention with frequency-differential networks to incorporate medical topological priors into implicit spatiotemporal dependency learning. Although these variants enhance time series modeling via attention-based techniques, their reliance on high-dimensional weight matrices to implicitly encode dependencies between time steps and variables hinders interpretability. This limitation is particularly severe in medical applications, where it undermines the trustworthiness of high-risk tasks such as disease diagnosis and risk prediction.
3 Preliminaries and Problem Formulation
We define the MedTS classification task as a supervised learning problem over time series data with hierarchical subject-level structure. The dataset is represented as , where is the -th sample with time steps and channels, is a multi-label vector over categories, and is the subject ID. Samples from the same subject share the same label, i.e., , where represents the set of samples from subject . We consider two dataset partitioning strategies: (i) Subject-Dependent Split, where the dataset is split at the sample level such that samples from the same subject may appear in both training and test sets, which can cause information leakage and yield overly optimistic performance estimates; and (ii) Subject-Independent Split, where the dataset is split at the subject level, ensuring that each subject appears only in one of the train, validation, or test sets, simulating real-world diagnostic scenarios but introducing challenges due to inter-subject variability. The goal is to train an encoder to obtain a latent representation , which is then passed to a classifier to predict the label .
@@ 4.1 Channel-Imposed Fusion
In this section, we analyze the data fusion process using Singular Value Decomposition (SVD) Golub and Reinsch (1971); Klema and Laub (1980). Suppose the original time-series data matrix has dimensions , where represents the temporal length and denotes the number of channels. We partition along the channel dimension into two submatrices: the first channels form , while the last channels form . The goal of fusion is to combine these two submatrices into a new representation:
@@
Gradient Coupling Mechanism: The coefficients and are optimized through gradient descent, where their updates are coupled due to shared dependencies on and . The gradients as:
Specifically, the gradient computations of the loss function with respect to and include information from each other’s signals, requiring their updates to simultaneously consider the influence of both forward and backward channels. This coupling mechanism enables the model to learn cross-channel interaction patterns, enhancing feature representation and generalization capability.
In Appendix B, we visualize scenarios with highly and generally correlated channel information.
@@ 4.2 HM-BiTCN Structure Design and Theoretical Analysis
Modeling short-term and long-term dependencies in time series data is challenging. Traditional CNNs excel at capturing local features but have limited receptive fields, hindering long-range dependency learning. Transformer-based methods effectively model long-term dependencies, but their complex design lacks interpretability, which is a key issue in medical time-series classification.To address these limitations, TCNs use causal convolutions for explicit temporal modeling and dilated convolutions to expand the receptive field, overcoming the constraints of traditional CNNs. Building on the advantages of TCN, we propose the HM-BiTCN, which combines the benefits of dilated convolutions, bidirectional causal convolution, and residual connections. This approach allows for better capture of temporal dependencies while preserving causality. Our goal is to simplify the model structure through explicit design principles while improving its interpretability.
4.2.1 Dilated Convolution
Dilated convolution expands the receptive field without significantly increasing computational cost Yu and Koltun (2015). For a 1D input sequence , its output is defined as , where is the current time step, is the kernel size, is the dilation factor, and is the weight at the -th position in the kernel. Increasing effectively enlarges the receptive field, enabling the network to capture longer-term temporal dependencies. When stacking multiple dilated convolutional layers, the receptive field grows progressively. For the -th layer, the receptive field can be expressed as , where is the dilation factor of the -th layer. By gradually increasing , the network captures temporal dependencies across both global and local scales, offering an effective way to model long-term dependencies in time series.
4.2.2 Bidirectional Causal Convolution Structure
4.2.3 Multi-Scale Feature Learning and Residual Connections
To further improve the model’s capacity to capture dependencies at different temporal scales, HM-BiTCN incorporates multi-scale feature learning and residual connections. Multi-scale Feature Learning: In HM-BiTCN, we employ a hierarchy of dilation factors that decrease layer by layer to capture temporal dependencies at multiple scales. Lower layers use larger dilation factors to expand the receptive field, aggregating long-range information and smoothing short-term noise in highly redundant medical time series; higher layers use smaller dilation factors to focus on local dependencies and capture fine-grained features. This coarse-to-fine, global-to-local design enables the network to extract broad patterns in its initial layers and refine precise details in its later layers, thereby enhancing adaptability across a wide range of time series tasks. Residual connections: Residual connections He et al. (2016) are introduced between the dilated convolutional layers to facilitate the efficient flow of information through the network. The residual connection is defined as , where is the convolutional output, and is the input. This design alleviates the vanishing gradient problem and improves the overall stability of the network during training.
@@ 5 Experiments
Datasets. (1) APAVA Escudero et al. (2006) is an EEG dataset where each sample is assigned a binary label indicating whether the subject has Alzheimer’s disease. (2) TDBrain van Dijk et al. (2022) is an EEG dataset with a binary label assigned to each sample, indicating whether the subject has Parkinson’s disease. (3) ADFTD Miltiadous et al. (2023b, a) is an EEG dataset with a three-class label for each sample, categorizing the subject as Healthy, having Frontotemporal Dementia, or Alzheimer’s disease. (4) PTB PhysioBank (2000) is an ECG dataset where each sample is labeled with a binary indicator of Myocardial Infarction. (5) PTB-XL Wagner et al. (2020) is an ECG dataset with a five-class label for each sample, representing various heart conditions. Table 4 provides information on the processed datasets.The data processing methodology is the same as that of Medformer Wang et al. (2024) and MedGNN Fan et al. (2025).
Baselines. We compare with 12 state-of-the-art time series transformer methods: Autoformer Wu et al. (2021), Crossformer Zhang and Yan (2022), FEDformer Zhou et al. (2022), Informer Zhou et al. (2021), iTransformer Liu et al. (2024b), MTST Zhang et al. (2024), Nonformer Liu et al. (2022), PatchTST Nie et al. (2023), Reformer Kitaev et al. (2019), vanilla Transformer Vaswani et al. (2017), Medformer Wang et al. (2024), MedGNN Fan et al. (2025).
Implementation. We employ six evaluation metrics: accuracy, precision (macro-averaged), recall (macro-averaged), F1 score (macro-averaged), AUROC (macro-averaged), and AUPRC (macro-averaged). The training process is conducted with five random seeds (41-45) on fixed training, validation, and test sets to compute the mean and standard deviation of the models. All experiments were conducted using an NVIDIA RTX 3090 GPU and implemented with PyTorch version 1.11.0 Paszke et al. (2017).

### v2

@@ Abstract
Medical time series (MedTS) such as EEG and ECG are critical for clinical diagnosis, yet existing deep learning approaches often struggle with two key challenges: the misalignment between domain-specific physiological knowledge and generic architectures, and the inherent low signal-to-noise ratio (SNR) of MedTS. To address these limitations, we shift from a conventional model-centric paradigm toward a data-centric perspective grounded in physiological principles. We propose Channel-Imposed Fusion (CIF), a method that explicitly encodes causal inter-channel relationships by linearly combining signals under domain-informed constraints, thereby enabling interpretable signal enhancement and noise suppression. To further demonstrate the effectiveness of data-centric design, we develop a simple yet powerful model, Hidden-layer Mixed Bidirectional Temporal Convolutional Network (HM-BiTCN), which, when combined with CIF, consistently outperforms Transformer-based approaches on multiple MedTS benchmarks and achieves new state-of-the-art performance on general time series classification datasets. Moreover, CIF is architecture-agnostic and can be seamlessly integrated into mainstream models such as Transformers, enhancing their adaptability to medical scenarios. Our work highlights the necessity of rethinking MedTS classification from a data-centric perspective and establishes a transferable framework for bridging physiological priors with modern deep learning architectures. The complete source code supporting this study is publicly available at the following Link: https://github.com/Xi-Mu-Yu/CIF.
@@ 1 Introduction
Medical time series (MedTS) data, such as electroencephalogram (EEG) and electrocardiogram (ECG) signals, are widely used in clinical settings to monitor patient health and play a crucial role in diagnosing neurological and cardiovascular diseases Arif et al. (2024); Xiao et al. (2023); Zhu et al. (2025); Wang et al. (2024b; 2025b). Accurate classification of these signals enables early anomaly detection, personalized treatment, and optimized therapy planning, ultimately improving patient outcomes and healthcare efficiency Liu et al. (2024a); Tian et al. (2023). With advances in deep learning, CNN-based models like EEGNet Lawhern et al. (2018) can automatically extract informative features from raw signals, significantly improving classification performance.
In recent years, Transformer models Vaswani et al. (2017), originally inspired by the self-attention mechanism Bahdanau et al. (2014), have achieved remarkable progress in time series modeling, particularly in capturing long-range dependencies and global contextual information Liu et al. (2021); Zhou et al. (2021). By mapping sequential data into high-dimensional token embeddings, Transformers are able to implicitly model complex temporal dependencies. Despite their success across a wide range of time series tasks, applying Transformer architectures to MedTS classification still faces several challenges, which can be summarized as follows: (1) Misalignment between domain-specific knowledge and generic architectures. Mainstream time series models, such as Autoformer Wu et al. (2021), Crossformer Zhang & Yan (2022), and Reformer Kitaev et al. (2019), have demonstrated strong performance in general domains such as weather forecasting and finance. However, as illustrated in Fig. 1, these approaches fail to achieve comparable effectiveness in MedTS classification tasks. This raises the urgent question of how to enhance the applicability of general-purpose models in medical scenarios. Moreover, MedTS often encode critical physiological characteristics—for example, conduction delays across ECG leads Auricchio et al. (2014) and rhythmic synchrony in EEG signals Palva & Palva (2014); Fries (2015)—which inherently reflect channel-level relationships. Unfortunately, such physiological dependencies are rarely considered in generic time series modeling frameworks. (2) Overemphasis on model optimization while neglecting the intrinsic low SNR of MedTS. Unlike general-purpose time series tasks, MedTS are characterized by pronounced low-SNR conditions Del Rio et al. (2011); Sraitih et al. (2022); Sharma (2017); Mohd Apandi et al. (2020); Jia et al. (2024), where noise and artifacts can easily overshadow critical physiological features. In such conditions, complex Transformer architectures do not always succeed in stably extracting effective representations, while simpler models (e.g., TCNs Bai et al. (2018)) may also experience more severe performance degradation. Indeed, recent Transformer-based methods tailored for MedTS, such as MedGNN Fan et al. (2025) and Medformer Wang et al. (2024a), primarily rely on architectural innovations, yet they fall short in fundamentally addressing the low-SNR challenge. This raises a key question: should breakthroughs in MedTS classification come from increasingly complex architectures, or from more principled data processing and representation strategies?
To address the aforementioned limitations, we depart from the traditional model-centric paradigm that relies on increasingly complex architectures to capture temporal dependencies, and instead propose a data-centric approach grounded in the physiological properties of medical time series. Following this principle, we introduce the Channel-Imposed Fusion (CIF) method, which explicitly encodes prior causal structures into feature representations. Specifically, CIF constructs new features through a linear combination of signals from different channels:
where and denote signals from two distinct channels, and and are coefficients predefined based on domain knowledge. When and take fixed values, they are not learned directly from patient data, but instead derived from two domain-specific prior hypotheses: (1) Physiological Coupling Hypothesis. For ECG signals, when two leads are highly correlated (e.g., P-wave polarity and morphology are consistent Platonov (2012)), setting achieves in-phase summation, thereby enhancing target signal components and improving the SNR. (2) Noise Suppression Hypothesis. In EEG recordings, ocular artifacts such as blinks often appear highly correlated in frontal electrodes Fp1 and Fp2 Croft & Barry (2000). To suppress such noise, we set , applying a differential fusion strategy to cancel common-mode interference. Here, the coefficients and serve as symbolic encodings of interpretable physiological principles, rather than exact data-driven estimates. When treated as learnable parameters, they can be fine-tuned under symbolic constraints imposed by prior knowledge (e.g., enforcing under coupling, and under noise suppression). This design maintains the interpretability of directional relationships (e.g., signal enhancement or cancellation) while allowing the model to adaptively adjust the magnitude of each coefficient based on the training data.
To emphasize the importance of data-centric approaches, we deliberately designed a simple yet effective model—the Hidden-layer Mixed Bidirectional Temporal Convolutional Network (HM-BiTCN)—to demonstrate that excellent performance does not necessarily require model complexity. The combination of CIF and HM-BiTCN not only outperforms Transformer-based methods on multiple medical datasets but also achieves new state-of-the-art (SOTA) results on general time series classification benchmarks. More importantly, the CIF method is not limited to the HM-BiTCN architecture itself; it exhibits strong transferability and can be seamlessly integrated into existing Transformer architectures, enhancing their adaptability to MedTS data. Our main contributions are:
Proposal of Channel-Imposed Fusion (CIF). We introduce CIF to explicitly model inter-channel relationships in medical time series, particularly suitable for signals with well-defined physiological structures such as EEG and ECG.
Design of HM-BiTCN based on CIF. By integrating CIF into HM-BiTCN, our method consistently outperforms existing SOTA models across multiple publicly available medical and non-medical time series classification datasets.
Methodological transferability. CIF is architecture-agnostic and can be seamlessly integrated into mainstream models such as Transformers, compensating for the limitations of traditional positional encodings in modeling channel-level correlations, and highlighting the paradigm shift from a model-centric to a data-centric perspective.
@@ 2 Related Work
Medical Time Series Classification. Medical time series analysis diverges fundamentally from general time series forecasting Wu et al. (2022a); Lu et al. (2024) by prioritizing pathological signature decoding over temporal extrapolation, with modalities like EEG Tang et al. (2021); Yang et al. (2023); Qu et al. (2020), ECG Xiao et al. (2023); Wang et al. (2023); Kiyasseh et al. (2021), and EMG [Xiong et al. (2021); Dai et al. (2022)] encoding distinct clinical semantics. Early methods were dominated by compact CNNs such as EEGNet Lawhern et al. (2018), which employs depthwise separable convolutions to efficiently extract spatio–temporal features while providing preliminary interpretability via feature‐map visualization. Subsequently, temporal convolutional networks (TCNs) Bai et al. (2018); Lin et al. (2019) leveraging dilated causal convolutions achieved parallelizable computation and extended receptive fields, surpassing LSTM‐based approaches Zhou et al. (2016); Shen & Lee (2016); Hochreiter & Schmidhuber (1997) on multiple medical signal classification benchmarks. Hybrid architectures such as EEG‐Conformer Song et al. (2022) combined convolutional front‐ends with Transformer self‐attention to capture both local and global dependencies and enabled attention‐based interpretability. More recently, fine-grained Transformer models such as Medformer Wang et al. (2024a) introduced cross-channel tokenization and dual-stage self-attention, setting new SOTA accuracy on several public datasets. The latest MedGNN Fan et al. (2025) further augments attention mechanisms with multi‐resolution graph learning to jointly model spatial multi‐scale channel dependencies and temporal dynamics.
@@
Model-centric Transformer-based time series methods. In time series analysis, Transformer-based models learn complex dependencies through diversified architectural designs: the vanilla Transformer Vaswani et al. (2017) first introduced multi-head self-attention and sinusoidal positional encoding to model temporal correlations globally; Informer Zhou et al. (2021) employs ProbSparse attention to select key time steps and compress sequence length, thereby reducing the computational cost of long-range dependencies; Reformer Kitaev et al. (2019) incorporates Locality-Sensitive Hashing (LSH) to reduce attention complexity to , making it suitable for ultra-long sequences; Autoformer Wu et al. (2021) proposes an Auto-Correlation mechanism that aggregates periodic subsequences to enhance the implicit capture of cyclic patterns; FEDformer Zhou et al. (2022) performs seasonal–trend decomposition in the frequency domain and uses compressed Fourier coefficients to enable cross-frequency attention interactions; Crossformer Zhang & Yan (2022) designs a two-stage attention mechanism across time and feature dimensions to implicitly fuse multivariate spatiotemporal couplings; iTransformer Liu et al. (2024b) innovatively treats time steps as channel dimensions and applies standard attention to implicitly learn nonlinear inter-variable relationships; PatchTST Nie et al. (2023) segments continuous time steps into patch-based tokens and uses a combination of local and global attention to capture multi-scale temporal patterns; Medformer Wang et al. (2024a) introduces multi-granularity patch embeddings and cross-channel attention for medical signals, implicitly modeling the heterogeneous couplings of physiological metrics; and MedGNN Fan et al. (2025) combines graph attention with frequency-differential networks to incorporate medical topological priors into implicit spatiotemporal dependency learning.
@@ 3.1 Channel-Imposed Fusion
In this section, we analyze the data fusion process using Singular Value Decomposition (SVD) Golub & Reinsch (1971); Klema & Laub (1980); Harner (1990); Lagerlund et al. (1997). Suppose the original time-series data matrix has dimensions , where represents the temporal length and denotes the number of channels. As shown in Fig.2, we partition along the channel dimension into two submatrices: the first channels form , while the last channels form . The goal of fusion is to combine these two submatrices into a new representation:
@@
It is worth noting that no matter how the channels are selected for partitioning and fusion, this operation is meaningful. Each group of channels contains a portion of the time-series information, and by performing SVD decomposition and linear combination on different channel subsets, one can explore their potential similarities and complementarities, thereby achieving either redundancy reduction or diversity enhancement. Therefore, this partitioning strategy is generally applicable and is not limited to the first and the last channels.
@@ 3.2 HM-BiTCN Structure Design and Theoretical Analysis
To demonstrate the value of data-centric approaches and to show that simple models can also achieve strong performance, as shown in Figure3 (a), the conventional TCN only considers unidirectional causal relationships in time series. In contrast, as shown in Figure3 (c), our proposed HM-BiTCN simply introduces bidirectional relationships and performs feature mixing at each layer. This design not only further enhances the signals with improved SNR through CIF but also captures both forward and backward causal structures, thereby making fuller use of the information contained in the data. See Appendix B and Appendix C for details. We must emphasize that such a simple modification does not constitute an innovation in itself. Our motivation is to demonstrate that simple models, when combined with a data-centric approach, can also surpass existing SOTA models.
@@ 4 Experiments
Medical Time Series Datasets. (1) APAVA Escudero et al. (2006) is an EEG dataset where each sample is assigned a binary label indicating whether the subject has Alzheimer’s disease. (2) TDBrain van Dijk et al. (2022) is an EEG dataset with a binary label assigned to each sample, indicating whether the subject has Parkinson’s disease. (3) ADFTD Miltiadous et al. (2023b; a) is an EEG dataset with a three-class label for each sample, categorizing the subject as Healthy, having Frontotemporal Dementia, or Alzheimer’s disease. (4) PTB PhysioBank (2000) is an ECG dataset where each sample is labeled with a binary indicator of Myocardial Infarction. (5) PTB-XL Wagner et al. (2020) is an ECG dataset with a five-class label for each sample, representing various heart conditions.
Baselines. We compare with 12 state-of-the-art time series transformer methods: Autoformer Wu et al. (2021), Crossformer Zhang & Yan (2022), FEDformer Zhou et al. (2022), Informer Zhou et al. (2021), iTransformer Liu et al. (2024b), MTST Zhang et al. (2024), Nonformer Liu et al. (2022), PatchTST Nie et al. (2023), Reformer Kitaev et al. (2019), vanilla Transformer Vaswani et al. (2017), Medformer Wang et al. (2024a), MedGNN Fan et al. (2025).
Implementation. We employ six evaluation metrics: accuracy, precision (macro-averaged), recall (macro-averaged), F1 score (macro-averaged), AUROC (macro-averaged), and AUPRC (macro-averaged). The training process is conducted with five random seeds (41-45) on fixed training, validation, and test sets to compute the mean and standard deviation of the models. All experiments were conducted using an NVIDIA RTX 3090 GPU and implemented with PyTorch version 1.11.0 Paszke et al. (2017). We consider two dataset partitioning strategies: (i) Subject-Dependent Split, where the dataset is split at the sample level such that samples from the same subject may appear in both training and test sets, which can cause information leakage and yield overly optimistic performance estimates; and (ii) Subject-Independent Split, where the dataset is split at the subject level, ensuring that each subject appears only in one of the train, validation, or test sets, simulating real-world diagnostic scenarios but introducing challenges due to inter-subject variability.